# A Taxonomy of Multi-Layered Runtime Guardrails for Designing Foundation Model-Based Agents: Swiss Cheese Model for AI Safety by Design

### v2

@@ IV-A1 Accuracy
Accuracy in FM-based agents is a critical concern, particularly regarding the generation of hallucinations, misinformation, and disinformation [45]. Hallucinations occur when models generate information that is factually incorrect or fabricated. Such inaccuracies can mislead users and damage the credibility of the systems [10]. Misinformation refers to the unintentional spread of false information, while disinformation involves the deliberate dissemination of falsehoods to deceive users [2]. These phenomena pose significant risks, as they can propagate false narratives and influence public opinion [20]. Guardrails mitigate these risks by detecting and filtering potentially inaccurate information [16]. For example, OpenAI uses guardrails to clearly label AI-generated content to prevent deepfakes and misinformation [46], such as in one case where it was reported to prevent misleading voters in the upcoming US elections [47, 48]. Therefore, guardrails improve the overall integrity of the systems [46, 1].
IV-A2 Privacy
In FM-based agents, privacy is a critical concern due to the risks of handling large amounts of personal and sensitive data. Without proper guardrails, these systems can inadvertently expose or misuse this data [12, 1]. One key privacy risk is data leakage, where sensitive information is inadvertently revealed through model outputs. For instance, an FM-based LLM trained on private conversations might accidentally generate text that includes personal information such as names, addresses, or financial details [49]. This leakage can occur through direct responses or statistical inferences. For example, an FM-based customer service chatbot might inadvertently include sensitive information. In April-May 2023, a notable incident involved Samsung employees leaking proprietary information into ChatGPT, leading to Samsung banning ChatGPT [50].
IV-A3 Security
Security in FM-based agents involves protecting them from malicious activities that could compromise their integrity and functionality [19, 14, 6]. For example, an FM-based system could be targeted by hackers to manipulate data, producing incorrect or harmful outputs that affect decision-making processes [51]. An incident similar to this reported in [52], where malicious users manipulated Microsoft’s Tay chatbot to produce inappropriate and offensive content, resulting in its shutdown. Additionally, FM-based agents are vulnerable to various threats [53]. Hackers can exploit vulnerabilities to access sensitive data, leading to breaches of confidentiality. Even with authorized access, there is a risk of data misuse by third-party providers [54]. FM-based agents are also prone to adversarial attacks, where specially designed queries extract sensitive information.
Guardrails mitigate these risks by detecting and responding to real-time threats, safeguarding system integrity and performance [51, 1]. They also prevent unauthorized access and the exploitation of vulnerabilities. For example, a financial FM-based system with guardrails can detect unusual transaction patterns and initiate defensive measures, such as temporarily suspending transactions [51]. Thus, guardrails better maintain confidentiality [55, 56, 49], system integrity [19, 10], and availability [57, 53, 18, 26, 11], prevent model misuse [54, 26].
IV-A4 Safety
FM-based agents face significant safety issues, particularly in generating harmful or misleading outputs. These issues can arise when models produce content that is inappropriate, offensive, or incorrect [3, 1]. For example, in applications where FM-based agents handle critical data, such as in medical diagnosis or or self-driving cars, incorrect outputs can lead to dangerous consequences [32]. Additionally, there is a risk of generating ethically or morally questionable content, which can damage the credibility and acceptance of the system [53].
Guardrails mitigate these risks by detecting and filtering out harmful output in real-time. For instance, in an FM-based virtual assistant used for mental health support, guardrails ensure the responses are benign and non-triggering. They also effectively detect and correct errors in real time, thus mitigating the effects of adversarial attacks [58, 59, 15, 60, 13].
IV-A5 Fairness
FM-based agents can face bias and discrimination in model outputs. These biases can emerge from the training data, model algorithms, or deployment context [61, 2]. For instance, a language model used in recruitment for initial screening of CVs and responses to selection criteria might inadvertently favor candidates from certain demographics, cultures, and languages [26, 8], which ultimately undermines the credibility of the systems. Guardrails address these fairness issues and ensure equitable outcomes through model parameter adjustments. For instance, in our recruitment example, they can analyze the model’s decisions and correct potential biases. Furthermore, guardrails promote transparency and accountability by providing insights into how decisions are made [18, 1], and avoid harmful stereotypes [53].
IV-A6 IP Protection
In FM-based agents, another key problem is the unauthorized use of generated content, where users might exploit outputs created by these systems without proper attribution or licensing, making them vulnerable to unauthorized use, duplication, or distribution [26, 55, 10]. Guardrails help mitigate these problems and better protect content’s copyright [49]. They also aid in compliance with copyright laws and licensing agreements to mitigate legal risks [19, 12]. Techniques such as watermarking, fingerprinting, and labeling ensure the copyright and ownership of data [10, 1].
IV-A7 Cost
@@
FM-based agents can incur significant costs due to errors, inefficiencies, or non-compliance with regulations. Without proper guardrails, agents might generate outputs that lead to financial losses, legal penalties, or damage to reputation [26, 12]. For example, an agent that provides incorrect financial advice could result in monetary losses for users and potential lawsuits against the provider. Implementing guardrails helps mitigate these risks by preventing erroneous or harmful outputs, thereby reducing potential costs associated with rectifying mistakes, handling legal issues, or rebuilding trust [1, 62]. Additionally, guardrails can optimize resource utilization by preventing unnecessary computations or actions, leading to cost savings in system operations [53].
IV-A8 Performance
Performance is a critical concern in FM-based agents, as they are expected to provide timely and efficient responses [24]. Without appropriate guardrails, agents might engage in resource-intensive computations that degrade system performance or lead to unacceptable response times [54]. Guardrails help manage performance by monitoring and controlling the computational resources used by the agent [64, 53]. For instance, they can limit the complexity of tasks the agent undertakes or prevent it from entering endless loops, ensuring consistent and efficient operation [49]. Moreover, guardrails can enhance performance by filtering out irrelevant or low-quality inputs, allowing the agent to focus on processing meaningful data [16, 1].
IV-B Internal Quality Attributes for Multilayer Runtime Guardrails (RQ2)
We identified eight key quality attributes essential for the design of runtime guardrails in FM-based agents. Below, we discuss these attributes in detail, highlighting their roles and importance.
IV-B1 Accuracy
Guardrail accuracy refers to the precision in identifying and mitigating risks, i.e., preventing undesired behaviors or outputs [45]. Functional accuracy emphasizes how well guardrails perform their intended tasks to meet user expectations and requirements. For instance, a guardrail used for a customer service chatbot must accurately block sensitive information from being disclosed. If a customer asks, ‘What is my current account balance?’ the guardrail must prevent disclosing the balance, instead providing a safe response like, ‘For security reasons, please check your balance through our secure app.’ If it fails to do so (false negatives) or incorrectly flags benign information (false positives), it compromises functional accuracy. Predictive accuracy involves the chatbot’s correct understanding and response to inquiries. We found that the accuracy of guardrails is evaluated using automated methods, primarily considering the systems’ performance [16], output quality [55], overall trustworthiness [10, 49], and system adaptability [26, 49].
IV-B2 Generalizability
@@
Generalizability in guardrails refers to their ability to function effectively across diverse applications [63]. Such guardrails ensure that the protective measures are not overly specific to a single use case but can adapt to various contexts and still perform reliably. For example, a chatbot must comply with different financial regulations across countries to ensure compliance without reconfiguration. Dong et al.[15] and Wang et al.[64] emphasize the need for guardrails that can extend their applicability to new domains without significant reconfiguration or loss of performance, even during unexpected inputs or data types. Furthermore, generalizability enhances the ability to handle diverse linguistic, cultural, and operational contexts to provide robust protection and to enhance the system’s resilience and reliability [12, 1].
IV-B3 Customizability
@@
Customizable guardrails provide tailored protection that meets specific requirements and supports diverse operational needs in FM-based agents [1, 65]. They allow for adjustments and configurations that align with particular operational goals, data characteristics, and regulatory environments. For example, a customer service chatbot can enable priorities for different guardrails and adjust data handling based on the user’s location, ensuring compliance with GDPR in Europe and CCPA in the U.S. [66]. Customizability also includes the capability to integrate user-defined rules and policies, ensuring that the guardrails can support diverse operational contexts and evolving needs [64].
IV-B4 Adaptability
@@
Adaptability in guardrails is known as their capability to adjust and remain effective under varying conditions and data landscapes as context evolves [26, 24]. This attribute ensures robust and continuous protection by dynamically responding to changes in input data, usage patterns, and emerging threats without manual reconfiguration [15]. For example, a customer service chatbot can automatically update its guardrails to detect and block new offensive terms during interactions. This includes the ability to incorporate new knowledge and advancements in threat detection techniques [54, 1].
IV-B5 Traceability
@@
The traceability attribute of guardrails tracks and records the origins, processes, and decision paths, such as input and output of FMs, external tools, etc. [27]. It involves maintaining detailed logs and records that can be audited to understand how decisions are made. For example, in a customer service chatbot, traceability ensures that every recommendation can be traced back to the data sources and algorithms used. This includes logging which data inputs influenced the response and what external APIs or tools were accessed, providing a clear audit trail for transparency and accountability. Traceability also aids in identifying the root causes of issues. This enables timely and accurate troubleshooting and improvement [26], and helps in maintaining user trust and meeting regulatory requirements [16, 10]. Additionally, comprehensive documentation of data sources and model modifications better support effective auditing and compliance checking [12].
IV-B6 Portability
@@
Portability in guardrails for FM-based agents refers to the ability of these protective measures to be easily adapted and applied across different foundation model (FM) systems and platforms [27]. This includes ensuring that the guardrails function consistently across various FM architectures and environments, thereby maintaining their effectiveness and integrity regardless of the underlying system [26]. For example, the same guardrail can be applied for content moderation in both a customer service chatbot and a social media platform, regardless of their underlying technology. The benefits of designing portable guardrails include compatibility across multiple programming languages and frameworks facilitate their integration into diverse technological stacks [49]. These capabilities ensure that the guardrails remain effective and operational as the system evolves or migrates to new environments. Portable guardrails also support seamless updates and improve scalability to maintain high standards of security and compliance while adapting to new technological advancements within systems [16].
IV-B7 Interoperability
@@
Interoperable guardrails work seamlessly across different systems and technologies [27]. They ensure that security, privacy, and compliance protocols can be applied consistently, even in heterogeneous environments that utilize varied software and hardware components, or diverse technological ecosystems [16, 67]. Guardrails that interface with various APIs and data formats also enable smooth communication and operation across different systems [26]. For example, they enable a customer service chatbot and internal support system to share data securely and consistently. This promotes cohesive and unified security management, reducing the complexity of maintaining multiple disparate protective measures [1], and better support collaborative efforts and data sharing [49].
IV-B8 Interpretability
@@
Interpretability refers to the clarity and transparency with which guardrails and protective measures operate. This allows users and stakeholders to understand how decisions are made and actions are taken by models. Thus increasing trust and accountability [10, 68]. For example, a chatbot in healthcare, can explain why certain advice is given or restricted. Transparent guardrails better facilitate auditing and compliance [18]. They also help users to understand that actions taken by guardrails can be clearly understood and verified [55]. This is essential for identifying and correcting errors, as well as for ensuring that the system’s operations align with standards.
IV-C Guardrails Design Options (RQ3)
@@
This section presents a structured taxonomy for designing multi-layered runtime guardrails for FM-based agents, focusing on identifying various design alternatives.
IV-C1 Guardrail Actions
@@
Guardrail actions are crucial for addressing the specific needs of FM-based agents. We have identified the following guardrail actions as key elements for FM-based agents:
@@
  • Flag: The flag action is used to mark specific inputs, outputs, or operations within FM-based agents [16]. For example, unusual transactions requested by the FM-based agent can be flagged for human review to ensure they comply with organizational policies [1, 30].
  • Modify: The modify action allows for the adjustment of inputs or outputs of various components in the FM-based agents to meet specific requirements or standards [9]. For example, the system can modify the user prompts by adding more context and examples, making it easier for the FM to accurately interpret the user’s intentions and provide more relevant responses.
  • Validate: The validate action checks inputs, outputs, and intermediate results against predefined criteria to ensure they meet specified requirements or standards [70, 26]. For example, the customer service chatbot could validate if the output generated by the FM-based system contains any sensitive company information.
  • Prioritize: The prioritize action enables the FM-based system to allocate resources and attention based on the importance of specific tasks, e.g., processing urgent user queries first [28].
  • Rate limit: The rate limit action controls the frequency and volume of requests or outputs processed by the system/component within a given time frame [16, 53]. For example, a rate limit can be set to restrict the number of API calls a single user can make.
  • Parallel calls: The parallel calls action can send multiple requests to the system/component to improve responsiveness, e.g., a user can send a prompt to the system multiple times at the same time and select the better response [16, 53].
@@
  • Fall back: When the system is unable to handle a request, the fall back action can redirect to the previous state or an alternative solution [13, 71, 16].
@@
  • Isolate: The isolate action involves segregating a specific entity (e.g., user) or component to prevent interaction with the system [57, 60, 19]. For example, a system might isolate a compromised narrow AI model suspected of being poisoned with malicious data in a sandbox environment, preventing potential harm to the main system.
  • Simulate: The simulate action involves running tests in a controlled environment to predict and analyze potential outcomes [1]. For instance, a system might simulate different configurations to determine most effective plan.
  • Redundancy: The redundancy action involves implementing backup processes or components to ensure continuity and reliability in case of failures [26, 16]. For example, an agent can implement two similar workflows in parallel, so if one workflow encounters an issue, the other can continue operating without interruption.
  • Log: The log action involves recording system activities, interactions, and events [70, 11]. For instance, logging all user interactions with an FM-based chatbot allows for the analysis of user interests.
TABLE V: A Mapping of Targets to Guardrail Actions
Type Targets Actions
FM-based agents Prompts Block, Filter, Flag, Modify, Validate, Prioritize, Rate Limit, Parallel Calls, Retry, Defer, Isolate, Log
Models Block, Filter, Flag, Modify, Validate, Prioritize, Rate Limit, Parallel Calls, Retry, Fall Back, Human Intervention, Defer, Isolate, Simulate, Redundancy, Log
External Data Block, Filter, Flag, Modify, Validate, Prioritize, Rate Limit, Parallel Calls, Retry, Defer, Isolate, Simulate, Redundancy, Log
Non-AI Components Block, Filter, Flag, Modify, Validate, Prioritize, Rate Limit, Parallel Calls, Retry, Fall Back, Human Intervention, Defer, Isolate, Simulate, Redundancy, Log
FM-based Agents Goals Block, Filter, Flag, Modify, Validate, Prioritize, Retry, Human Intervention, Defer, Simulate, Log
Context Block, Filter, Flag, Modify, Simulate, Log
Memory Block, Filter, Flag, Modify, Validate, Prioritize, Retry, Human Intervention, Isolate, Simulate, Log
Reasoning Flag, Modify, Validate, Human Intervention, Log
Plans Block, Flag, Modify, Validate, Retry, Fall Back, Human Intervention, Defer, Simulate, Log
Actions Block, Flag, Modify, Validate, Retry, Prioritize, Parallel Calls, Fall Back, Human Intervention, Defer, Simulate, Log
Intermediate Results Flag, Validate, Human Intervention, Log
Final Results Block, Filter, Flag, Modify, Validate, Retry, Fall Back, Human Intervention, Log
Tools Block, Prioritize, Rate Limit, Parallel Calls, Retry, Fall Back, Human Intervention, Defer, Log
Other Agents Block, Flag, Rate Limit, Parallel Calls, Retry, Fall Back, Human Intervention, Defer, Isolate, Simulate, Log
Intermediate Results Flag, Validate, Human Intervention, Log
Final Results Block, Filter, Flag, Modify, Validate, Retry, Fall Back, Human Intervention, Log
IV-C2 Targets for Guardrails
The key targets guardrails can be applied to include prompts, models, external data, non-AI components, and agent-specific targets. Table V provides an overview of targets and corresponding guardrail actions.

### v3

@@ IV-A1 Accuracy
Accuracy in FM-based agents is crucial, particularly in mitigating issues such as hallucinations, misinformation, and disinformation [45]. Hallucinations occur when models generate information that is factually incorrect. Such inaccuracies can mislead users and damage the credibility of the agent [10]. Misinformation refers to the unintentional spread of false information, while disinformation involves the deliberate dissemination of falsehoods to receive users [20]. For example, OpenAI uses guardrails to clearly label AI-generated content to prevent deepfakes and misinformation [46]. One such case has been reported to prevent misleading voters in last US elections [47, 48].
IV-A2 Efficiency
@@
Efficiency is crucial in FM-based agents, as users expect fast, efficient responses [24]. Without guardrails, agents risk engaging in resource-intensive tasks that slow down response times [54]. By dynamically managing resources across multiple layers, these guardrails prevent inefficiencies, such as endless loops, and filter irrelevant inputs, ensuring that agents focus on processing meaningful data [64, 16, 3]. Additionally, FM-based agents can incur significant costs due to errors, inefficiencies, or non-compliance with regulations. Without proper guardrails, agents might generate outputs that lead to financial losses, legal penalties, or damage to their reputation [26, 12]. For example, an agent that provides incorrect financial advice could result in monetary losses for users and potential lawsuits against the provider.
IV-A3 Privacy
Privacy in FM-based agents poses risks due to handling sensitive data, where data leakage might expose personal information [12, 1]. This leakage can occur through direct responses or statistical inferences, or inadvertent revelations through model outputs. In April-May 2023, a notable incident involved Samsung employees leaking proprietary information into ChatGPT, leading to Samsung banning ChatGPT [50].
IV-A4 Security
Security in FM-based agents involves protecting them from malicious activities that could compromise their integrity and functionality [19, 14, 6]. For example, an FM-based agent could be targeted by hackers to manipulate data, producing incorrect or harmful outputs that affect decision-making processes [51]. An incident reported in [52] described how malicious users manipulated Microsoft’s Tay chatbot to produce inappropriate (offensive) content, leading to its shutdown. FM-based agents are also vulnerable to hacks that may breach data confidentiality [53]. Even with authorized access, there is a risk of data misuse by third-party providers [54]. Moreover, FM-based agents are prone to adversarial attacks, where specially designed queries extract sensitive information. Guardrails mitigate these risks by detecting and responding to real-time threats across various operational layers, safeguarding agent integrity [19, 10], confidentiality [55, 56, 49], availability [57, 53, 18, 26, 11] and performance [51, 1].
IV-A5 Safety
FM-based agents face significant safety issues, particularly in generating harmful or misleading outputs. These issues can arise when models produce content that is inappropriate, offensive, or incorrect [3]. These issues are critical in contexts where FM-based agents handle critical data like medical diagnosis or self-driving cars, where inaccurate outputs could have severe consequences [32]. Additionally, there is a risk of generating questionable content, which can damage the credibility and acceptance of the agent [53].
IV-A6 Fairness
FM-based agents can face bias and discrimination in model outputs. These biases can emerge from the training data, model algorithms, or deployment context [61, 2]. For instance, an agent used in recruitment for screening CVs might inadvertently favor candidates from certain demographics, cultures, and languages [26, 8], affecting credibility.
IV-A7 Compliance
Compliance in FM-based agents involves adhering to legal and regulatory standards [16, 20]. These issues are critical because non-compliance can lead to legal penalties, reputational damage, and loss of user trust. Runtime guardrails reduce these risks by ensuring alignment with data protection regulations, industry standards, and guidelines through continuous monitoring at multiple levels [54, 26]. Additionally, these guardrails assist in automating compliance checks. They ensure that all aspects of the FM-based agent’s operations align with the necessary legal and regulatory frameworks [62, 36], and better support internal audits and external regulatory reviews [12]. For example, FM-based agents may unintentionally facilitate unauthorized use of generated content, making it vulnerable to duplication or improper distribution [26, 55, 10]. Guardrails operating in real time help mitigate these risks by detecting and restricting unauthorized access, ensuring better copyright protection [49]. Techniques such as watermarking, fingerprinting, and labeling are applied across different layers to ensure the ownership and compliance with licensing laws [10, 1].
IV-A8 Generalizability
@@
Generalizability in guardrails for FM-based agents refers to their ability to function effectively in real-time across multiple layers and diverse scenarios without prior configurations [63]. Such guardrails ensure that protective measures are not overly specific to a single use case but can adapt to various contexts and still perform reliably across layers. The agents’ ability to handle diverse linguistic, cultural, and operational contexts is essential to provide robust protection, resilience, and reliability and is ensured by the generalizability attribute [12, 1]. Guardrails that can extend their applicability to new domains without significant reconfiguration or degradation in performance, even during unexpected inputs or data types, are essential [15, 64].
IV-A9 Customizability
@@
Customizable guardrails provide tailored protection that meets specific requirements and supports diverse operational needs in FM-based agents [1, 65]. The multi-layered runtime approach allows for customization at different layers to enable fine-grained control over the agent’s behavior during execution, such as adjustments and configurations that align with particular operational goals, data characteristics, and regulatory environments. For example, a customer service chatbot can enable priorities for different guardrails and adjust data handling based on the user’s location and ensuring compliance with regulation.
IV-A10 Adaptability
@@
Adaptability in guardrails is known as their capability to adjust and remain effective under varying conditions and data landscapes as context evolves [26, 24]. This attribute ensures robust and continuous protection by dynamically responding to changes in input data, usage patterns, and emerging threats without manual reconfiguration [15]. For example, a customer service chatbot can automatically update its guardrails to detect and block new offensive terms during interactions. This includes incorporating new knowledge and advancements in threat detection techniques [54, 1].
IV-A11 Traceability
@@
The traceability attribute of guardrails tracks and records the origins, processes, and decision paths, such as input and output of FMs, external tools, etc. [27]. It involves maintaining detailed logs and records that can be audited to understand how decisions are made. For example, in a customer service chatbot, traceability ensures that every recommendation can be traced back to the data sources and algorithms used. This provides a clear audit trail for transparency and accountability. Traceability also aids in identifying the root causes of issues to enable timely and accurate troubleshooting and improvement [26], and helps in maintaining user trust and meeting regulatory requirements [16, 10]. Comprehensive documentation of data sources and model modifications also better support effective auditing and compliance checking [12].
IV-A12 Portability
@@
Portability in guardrails for FM-based agents refers to the ability of these protective measures to be easily adapted and applied across different FM-based agents [27]. Multiple layer runtime guardrails allow individual layers to be transferred and integrated into different agents with minimal adjustments in real time. This includes ensuring that they function consistently across various FM architectures and environments, thereby maintaining their effectiveness and integrity regardless of the underlying technologies [26]. For example, the same guardrail can be applied for content moderation in both a customer service chatbot and a social media platform, regardless of their underlying technology. The benefits of designing portable guardrails include compatibility across multiple programming languages and frameworks facilitate their integration into diverse technological stacks [49]. These capabilities ensure that the guardrails remain effective and operational as the agent evolves or migrates to new environments. Portable guardrails also support seamless updates and improve scalability to maintain high standards of security and compliance while adapting to new technological advancements within agents [16].
IV-A13 Interoperability
@@
Interoperable guardrails work seamlessly across differing agents, technologies and interface effectively with various components and services within different agents [27]. They ensure that security, privacy, and compliance protocols can be applied consistently, even in heterogeneous environments that utilize varied software and hardware components, or diverse technological ecosystems [16, 67]. Guardrails that interface with various APIs and data formats also enable smooth communication and operation across different agents [26]. For example, they enable a customer service copilot and internal support system to share data securely and consistently. This promotes cohesive and unified security management, reducing the complexity of maintaining multiple disparate protective measures [1], and better support collaborative efforts and data sharing [49].
IV-A14 Interpretability
@@
Interpretability refers to the clarity and transparency with which guardrails and protective measures operate. Interpretability allows better inspection and understanding of each layer’s function during execution. This allows users and stakeholders to understand how decisions are made and actions are taken by models. Thus increasing trust and accountability [10, 68]. For example, a chatbot in healthcare, can explain why certain advice is given or restricted. Transparent guardrails better facilitate auditing and compliance [18]. They also help users to understand that actions taken by guardrails can be clearly understood and verified [55]. This is essential for identifying and correcting errors, as well as for ensuring that the agent’s operations align with ethical and regulatory standards.
IV-B Design Options of Guardrails
@@
This section presents a structured taxonomy for designing guardrails, focusing on identifying various design alternatives.
IV-B1 Actions
@@
Guardrail actions are crucial for addressing the specific needs of FM-based agent artifacts. We have identified the following guardrail actions that can be applied to FM-based agents:
@@
  • Flag: The flag action is used to mark specific inputs, outputs, operations within FM-based agents [16]. For example, unusual transactions requested by the FM-based agent can be flagged for human review to ensure they comply with organizational policies [1, 30].
  • Modify: The modify action allows for the adjustment of inputs or outputs of various components in FM-based agents to meet specific requirements or standards [9]. For example, the user prompts can be modified by adding more context and examples, making it easier for the FM to accurately interpret the user’s intentions and provide more relevant responses.
  • Validate: The validate action checks agent artifacts against predefined criteria to ensure they meet specified requirements or standards [70, 26]. For example, the plan generated by FM-based agents should be validated, e.g., through external verifier [79], to ensure it is compliant with regulatory policies.
  • Parallel calls: The parallel calls action can send multiple requests to the agent/component to improve responsiveness, e.g., a user can send a prompt to the agent or an external service multiple times at the same time and select the better response [16, 53].
@@
  • Fall back: When one step in the workflow cannot be executed successfully, the fall back action redirect to the previous step and state [13, 71, 16].
@@
  • Isolate: The isolate action involves segregating a specific entity (e.g., user) or component to prevent interaction with the agent [57, 60, 19]. For example, an agent might isolate a compromised narrow AI model suspected of being poisoned with malicious data in a sandbox environment, preventing potential harm to the agent.
  • Redundancy: The redundancy action involves implementing backup processes or components to ensure continuity and reliability in case of failures [26, 16]. For example, two sensors can be deployed to detect context information for an agent.
  • Evaluate: The evaluate action involves assessing the results [1]. For instance, an agent might ask another agent to evaluate its intermediate or final results.
TABLE V: A Mapping of Agent Targets to Guardrail Actions
Type Targets Guardrail Actions
Pipeline Prompts Block, filter, flag, modify, parallel calls, retry, defer, evaluate
Intermediate results Flag, human intervention, evaluate
Final results Block, filter, flag, modify, retry, fall back, human intervention, evaluate
Artifacts Goals Validate, block, flag, modify, human intervention, defer
Context Block, filter, flag, modify, evaluate
Memory Block, filter, flag, modify, retry, human intervention, isolate, evaluate
Reasoning Flag, modify, validate, human intervention
Plans Block, flag, modify, validate, retry, fall back, human intervention, defer
Workflows Validate, parallel calls, retry, fall back, human intervention, defer, evaluate
Tools Block, parallel calls, retry, fall back, human intervention, defer, evaluate
Knowledge bases Block, filter, flag, modify, retry, isolate, evaluate, redundancy
Other agents Block, flag, parallel calls, retry, fall back, human intervention, defer, isolate, evaluate
FMs Block, filter, flag, modify, parallel calls, retry, fall back, human intervention, isolate, evaluate, redundancy
IV-B2 Targets
Guardrail actions can be applied to various targets across multi-layers, including both pipelines and artifacts. Some guardrails are applied the the entire pipeline (including prompts, intermediate results, and final results), while others are focused on specific artifacts (covering goals, context, reasoning, plans, memory, tools, knowledge bases, other agents, FMs). Table V provides an overview of agent targets and corresponding guardrail actions.